# Joint Denoising and Fusion with Short- and Long-exposure Raw Pairs

### v4

@@ 4.3 Loss Functions
5 Experiments
5.1 Experimental Setup
Datasets and Metrics. We utilize the proposed RealRaw-HDR dataset for training. For testing, we first evaluate our method on the dataset, which is synthesized using the data formation pipeline. This dataset was not utilized during training. This test set contains 30 sets of test samples containing different exposure ratios (i.e., 4, 8, and 16) with a resolution of $4096\times 2176$. To validate the generalization ability of our method on real data, we utilize a FUJI-FILM GFX50S II camera to capture seven sets of real-world bracketed exposure raw images and the corresponding static images for generating the ground truth. Furthermore, to validate the generalization ability of our method on HDR sensors, we utilize Chen’s test dataset [5] for cross-validation, which has short-long raw images captured by a Sony IMX267 image sensor.
We perform a quantitative evaluation using the PSNR, SSIM, and CIE L*a*b* space  [^3] CIE L*a*b* is a color space specified by the International Commission on Illumination. It describes all the colors visible to the human eye and was created to serve as a device-independent model to be used as a reference. [45, 13, 20] (also known as $\Delta E$). $\Delta E$ can effectively assess chromaticity, contrast, and color accuracy variations within HDR images rather than exclusively concentrating on luminance differences (HDR-VDP-2). It offers a comprehensive quality evaluation by measuring the disparity between two HDR images within the CIE Lab color space.
$\Delta E=\|\widetilde{I}^{out}_{lab}-{I}^{gt}_{lab}\|_{2},$ (11)
where $\widetilde{I}^{out}_{lab}$ and ${I}^{gt}_{lab}$ are the CIE Lab version of the predicted HDR image and ground truth, respectively.
@@ 5.1 Experimental Setup
Figure 8: Visual comparison of state-of-the-art HDR reconstruction methods on RealRaw-HDR dataset with realistic exposure ratios.
@@
5.2 Comparison with the State-of-the-art Methods
We choose several representative low-level vision methods for comparisons, including four HDR methods based on sRGB images (AHDRNet [37], DeepHDR [16], NHDRRNet [40], HDR-Transformer[22]), as well as two methods for denoising raw images (SGN [10] and UNet-SID [3]). For fair comparisons, we re-train all the methods using the RealRaw-HDR dataset. Additionally, for AHDRNet, DeepHDR, HDR-Transformer, and NHDRRNet, we modify the network inputs to accommodate dual-exposure raw images. Similarly, for SGN and UNet-SID, we concatenate the long- and short-exposure images as inputs.
@@ 5.2 Comparison with the State-of-the-art Methods
Evaluation on synthetic dataset. We first evaluated our method on a synthetic dataset generated using the raw HDR data formation pipeline. The quantitative comparison results are shown in Tab. 2. The results clearly show that our method outperforms previous methods in almost all metrics on the synthetic dataset. Notably, our lightweight-efficient RepUNet model has fewer parameters (0.82M) and remarkably low GFLOPs (127G FLOPs). This efficiency allows us to process two 4K Bayer raw images at only 2.9ms using an NVIDIA RTX 3090 GPU. In contrast, other models with comparable performance necessitate significantly longer processing times. RepUNet achieves comparable performance with AHDRNet with only $4.5\%$ of its computational complexity (127G vs. 2848G). Fig 6 visualizes that our method can effectively eliminate noise and ghosting artifacts in the reconstructed HDR. In comparison, DeepHDR [16], NHDRRNet [40], and SGN [10] exhibit numerous artifacts in the palm motion region. However, our proposed RepUNet can reconstruct HDR images without ghosting (see rows 2 and 4).
Evaluation on HDR sensor raw dataset. To validate the generalization ability of our method on the real-world HDR sensor, we utilize Chen’s test dataset [5] for cross-validation, which has raw images captured by a Sony IMX267 image sensor. Compared with previous methods, our method achieves state-of-the-art performance in terms of both visual quality and quantitative metrics. The visual results from tests on the HDR sensor raw dataset (as shown in Fig. 7) indicate that DeepHDR, NHDRRNet, and SGN show noticeable ghosting, with NHDRRNet also suffering from severe color casts. Furthermore, results in Tab. 3 reveal that compared to AHDRNet [37], our method yields an improvement of more than 0.35 dB and 0.05 gain in PSNR and $\Delta E$, respectively, for scenes with an exposure ratio of 8. On average, our method attains gains exceeding 0.1 dB and 0.15 in these metrics.
Evaluation on FUJI raw dataset. We then evaluate our method on the FUJI raw datasets, which are real-world bracketed exposure raw images captured by the FUJI-FILM GFX50S II camera. Fig. 8 compares results from two high dynamic range scenes, where our method achieves significantly better visualization. Our method can recover both fine details in overexposed regions and rich colors in underexposed areas without introducing artifacts (see rows 1 and 2). Compared to AHDRNet, our method can effectively remove noise and preserve the structure of dark regions. Notably, the alignment module in DeepHDR, AHDRNet, HDR-Transformer, and NHDRRNet requires many line buffers, making it challenging to deploy on resource-limited edge devices. Notably, HDR-Transformer fails to perform inference even on RTX 3090 devices. In contrast, our method can alleviate ghost artifacts without relying on any alignment module and addresses color cast issues in raw images, as evident in the visual results of Fig. 6, Fig. 7, and Fig. 8.
@@
Table 6: Reparameterization ablation results. The FLOPs and Run Times are measured on the raw image of 4K resolution.
Method FLOPs Params Run Times All-Exposure Ratio=4 Ratio=8 Ratio=16
[…]
DualUNet 93.26G 0.82M 3.0 ms 39.7941 2.1202 39.8026 2.1036 40.5092 1.9790 39.0705 2.2780
[…]
Despite the fact that all models are trained on the RealRaw-HDR dataset, which is synthesized using the data formation pipeline, they consistently excel on both the synthetic test set and the real-world dataset. Particularly noteworthy is the remarkable performance achieved on the test set comprised of raw images captured by the HDR sensor [5]. These results serve as solid evidence of the generalizability of our proposed RealRaw-HDR dataset and the HDR data formation pipeline. It underscores the robustness inherent in our approach and its capacity to tackle real-world challenges. This not only validates the efficacy of our methodology but also highlights its considerable potential for diverse real-world applications.
@@ 5.3 Ablation Study
Generalization of our LDR-HDR pair formation pipeline. Our raw LDR-HDR pair formation pipeline is proposed to generate paired raw LDR-HDR data but can also be adapted to generate paired sRGB HDR data. To demonstrate such generalization, we transform the collected RealRaw-HDR dataset with a fixed ISP pipeline into the sRGB color space, named the Raw2RGB-HDR dataset. For comparison, we train the sRGB HDR method AHDRNet [37] on our Raw2RGB-HDR dataset (240 pairs) and Kalantari’s dataset [16] (taking the first two exposures as input, 74 pairs of images), respectively. The test dataset is from Kalantari’s dataset. Results in Tab. 7 show that AHDRNet trained on our Raw2RGB-HDR dataset outperforms the one trained on Kalantari’s dataset by 2.86 dB in PSNR. The performance gains benefit from an efficient and user-friendly data acquisition pipeline that generates more trainable data pairs. The results demonstrate that our data pipeline is also effective in generating paired LDR-HDR data in sRGB space.
Table 7: We train the sRGB HDR method AHDRNet on our Raw2RGB-HDR dataset (240 pairs) and Kalantari’s dataset, respectively. RealRGB-HDR is obtained by processing the RealRaw-HDR dataset.
@@
Loss Functions. To test the effects of alignment-free and motion-aware short-exposure-first selection loss and Bayer loss, we set the $L2$ joint $L_{\textrm{ssim}}$ loss as the baseline loss and step-by-step modify the loss function combination. Tab. 4 and 9 show that adding the AMSS loss and Bayer loss steadily improves visual quality and quantitative results. RepUNet with joint loss achieves the best results, outperforming the baseline by 0.5 dB in PSNR and by 0.16 in $\Delta E$ on average. As Fig. 9 shows, alignment-free and motion-aware short-exposure-first selection loss (AMSS-Loss) effectively suppresses the ghosting artifacts (see columns 3 and 4). Meanwhile, our proposed Bayer loss can alleviate the color cast (see columns 2 and 3).
Figure 9: Visual results of RepUNet and its baseline variants. By combining these loss functions, our model can effectively produce top-notch results for both motion and saturated areas.
@@
6 Discussion and Future Work
Although our method can produce high-quality HDR images in real scenes, it has several limitations that may inspire future work.
Our data synthesis pipeline employs noise simulation to synthesize long and short-exposure images. However, we have found that if the noise distribution of the images we test is significantly different from the noise model we have measured, the results of the reconstruction are often unsatisfactory. Nevertheless, for practical engineering applications, it is important to accurately measure each type of image sensor noise model to achieve efficient raw denoising.
Furthermore, our work does not cover scenarios involving flicker. This is because the flicker problem generally arises in short-exposure images, and it is difficult to determine which region is better when flicker and motion occur simultaneously. Unfortunately, there is no clear-cut solution that can effectively address scenes where both flicker and motion are present. To overcome this challenge, it may be necessary to create a flicker model to generate a substantial amount of training data.
@@ 7 Conclusion
We propose a novel data formation pipeline for raw LDR-HDR pairs based on HDR sensor imaging systems and construct a new high-quality raw dataset and raw LDR-HDR pairs dataset. This dataset can efficiently support many kinds of dual-exposure HDR sensor development, which provides vital data support for realizing the new generation of AI ISPs. Meanwhile, we develop a lightweight and efficient RepUNet network that uses structural reparameterization for HDR image reconstruction. Our evaluation has shown that this model can be easily deployed on mobile devices. We propose a plug-and-play alignment-free and motion-aware short-exposure-first selection loss and Bayer loss to address the issue of ghosting artifacts and color cast. Extensive experiments demonstrate that our approach achieves state-of-the-art performance in both visual quality and quantitative metrics.

### v5

@@ 4.5 Loss Functions
5 Experiments
5.1 Experimental Setup
Datasets and metrics. We utilize the proposed RealRaw-HDR dataset for training. We first evaluate our method in the synthesized dataset. This test set contains 30 samples containing different exposure ratios (i.e., 4, 8, and 16) with a resolution of $4096\times 2176$. To validate the validity of our method on real data, we utilize a FUJI-FILM GFX50S II camera to capture seven sets of real-world bracketed exposure raw images and the corresponding static images for generating the ground truth. Furthermore, we also utilize the Chen [4] test dataset for cross-validation, which has short- and long-exposure raw pairs captured by a Sony IMX267 image sensor.
We perform a quantitative evaluation using the PSNR, SSIM, and CIE L*a*b* space  [^1] CIE L*a*b* is a color space specified by the International Commission on Illumination. It describes all the colors visible to the human eye and was created to serve as a device-independent model for reference. [43, 11, 19] (also known as $\Delta E$). $\Delta E$ can effectively assess chromaticity, contrast, and color accuracy variations within HDR images rather than exclusively concentrating on luminance differences (HDR-VDP-2). It offers a comprehensive quality evaluation by measuring the disparity between two HDR images within the CIE L*a*b* color space.
$\Delta E=\|\widetilde{I}^{out}_{lab}-{I}^{gt}_{lab}\|_{2},$ (20)
where $\widetilde{I}^{out}_{lab}$ and ${I}^{gt}_{lab}$ are the CIE L*a*b* version of the predicted HDR image and ground truth, respectively.
@@ 5.1 Experimental Setup
5.2 Comparison with the Other Methods
We choose several representative low-level vision methods for comparisons, including four HDR methods based on sRGB images (AHDRNet [34], DeepHDR [33], NHDRRNet [37], HDR-Transformer[21]), as well as two methods for denoising raw images (SGN [8] and UNet-SID [2]). For fair comparisons, we re-train all the methods using the RealRaw-HDR dataset. Additionally, for AHDRNet, DeepHDR, HDR-Transformer, and NHDRRNet, we modify the network inputs to accommodate dual-exposure raw images. Similarly, for SGN and UNet-SID, we concatenate the long- and short-exposure raw pairs as inputs.
Evaluation on synthetic dataset. We first evaluated our method on a synthetic dataset generated using the raw HDR data formation pipeline. The quantitative comparison results are shown in Tab. 2. The results clearly show that our method outperforms previous methods in almost all metrics on the synthetic dataset. Notably, our lightweight-efficient RepUNet model has fewer parameters (0.82M) and remarkably low GFLOPs (127G FLOPs). This efficiency allows us to process two 4K Bayer raw images at only 2.9ms using an NVIDIA RTX 3090 GPU. In contrast, other models with comparable performance necessitate significantly longer processing times. RepUNet achieves comparable performance with AHDRNet with only $4.5\%$ of its computational complexity (127G vs. 2848G). Fig 4 visualizes that our method can effectively eliminate noise and ghosting artifacts in the reconstructed HDR. In comparison, DeepHDR [33], NHDRRNet [37], and SGN [8] exhibit numerous artifacts in the palm motion region. However, our proposed RepUNet can reconstruct HDR images without ghosting (see rows 2 in Fig. 4).
Figure 7: Visual comparison of state-of-the-art HDR reconstruction methods on FUJI raw dataset.
@@ 5.2 Comparison with the Other Methods
Table 5: Reparameterization ablation results. The FLOPs and run times are measured on the raw image with a 4K resolution.
Method FLOPs Params Run Times All-Exposure Ratio=4 Ratio=8 Ratio=16
[…]
Base Model 93.26G 0.82M 3.0 ms 39.7941 2.1202 39.8026 2.1036 40.5092 1.9790 39.0705 2.2780
Evaluation on HDR sensor dataset. To validate the validity of our method on the real-world HDR sensor dataset, we utilize the Chen [4] test dataset for cross-validation, which has raw images captured by a Sony IMX267 image sensor. Compared with previous methods, our method achieves state-of-the-art performance in visual quality and quantitative metrics. The visual results from tests on the HDR sensor raw dataset (as shown in Fig. 6) indicate that DeepHDR, NHDRRNet, and SGN show noticeable ghosting, with NHDRRNet also suffering from severe color casts. Furthermore, results in Tab. 3 reveal that compared to AHDRNet [34], our method yields an improvement of more than 0.35 dB and 0.05 gain in PSNR and $\Delta E$, respectively, for scenes with an exposure ratio of 8. On average, our method attains gains exceeding 0.1 dB and 0.15 in PSNR and $\Delta E$.
@@
Evaluation on FUJI raw dataset. We then evaluate our method on the FUJI raw datasets, which are real-world bracketed exposure raw images captured by the FUJI-FILM GFX50S II camera. Fig. 7 compares results from two high dynamic range scenes, where our method achieves significantly better visualization. Our method can recover both fine details in overexposed regions and rich colors in underexposed areas without introducing artifacts (see rows 1 and 2). Compared to AHDRNet, our method can effectively remove noise and preserve the structure of dark regions. Notably, the alignment module in DeepHDR, AHDRNet, HDR-Transformer, and NHDRRNet requires many line buffers, making it challenging to deploy on resource-limited edge devices. HDR-Transformer fails to perform inference even on RTX 3090 devices. In contrast, our method can alleviate ghost artifacts without relying on any alignment module and addresses color cast issues in raw images, as evident in the visual results of Fig. 4, Fig. 6, and Fig. 7.
Although all models are trained on the RealRaw-HDR dataset, which is synthesized using the data formation pipeline, they consistently excel on both the synthetic test dataset and the real-world dataset. Particularly noteworthy is the remarkable performance achieved on the test dataset comprised of raw images captured by the HDR sensor [4]. These results are solid evidence of the generalizability of our proposed RealRaw-HDR dataset and the HDR data formation pipeline.
@@ 5.3 Ablation Study
Generalization of our LDR-HDR pair formation pipeline. Our raw LDR-HDR pair formation pipeline is proposed to generate paired raw LDR-HDR data but can also be adapted to generate paired sRGB HDR data. To demonstrate such generalization, we transform the collected RealRaw-HDR dataset with a fixed ISP pipeline into the sRGB color space, named the Raw2RGB-HDR dataset. For comparison, we train the sRGB HDR method AHDRNet [34] on our Raw2RGB-HDR dataset and Kalantari dataset [14] (taking the first two exposures as input, 74 pairs of images), respectively. The test dataset is from the Kalantari dataset. Results in Tab. 7 show that AHDRNet trained on our Raw2RGB-HDR dataset outperforms the one trained on the Kalantari dataset by 2.86 dB in PSNR. The performance gains benefit from an efficient and user-friendly data acquisition pipeline that generates more trainable data pairs. The results demonstrate that our data pipeline is also effective in generating paired LDR-HDR data in sRGB space.
Table 7: We train the sRGB HDR method AHDRNet on our Raw2RGB-HDR dataset and Kalantari dataset, respectively. RealRGB-HDR is obtained by processing the RealRaw-HDR dataset.
@@
Figure 8: Visual results of RepUNet and its baseline variants. Combining these loss functions allows our model to produce top-notch results for motion and saturated areas effectively.
Loss functions. To test the effects of alignment-free and motion-aware short-exposure-first selection loss and Bayer loss, we set the $L2$ joint $L_{\textrm{ssim}}$ loss as the baseline loss and step-by-step modify the loss function combination. Tab. 4 and Fig. 8 show that adding the AMSS and Bayer loss steadily improves visual quality and quantitative results. RepUNet with joint loss achieves the best results, outperforming the baseline by 0.5 dB in PSNR and by 0.16 in $\Delta E$ on average. As Fig. 8 shows, alignment-free and motion-aware short-exposure-first selection loss (AMSS-Loss) effectively suppresses the ghosting artifacts (see columns 3 and 4). Meanwhile, our proposed Bayer loss can alleviate the color cast (see columns 2 and 3).
@@ 6 Conclusion
In the paper, we proposed a Topological Convolution Block (TCB) for efficient and light-weight HDR design for mobile devices. Based on the proposed TCB, we further designed RepUNet, aiming at balancing hardware efficiency and PSNR/SSIM indexes. Furthermore, We propose a novel computational photography based pipeline for raw HDR image formation and construct a real-world raw HDR dataset – RealRaw-HDR. Meanwhile, we designed plug-and-play alignment-free and motion-aware short-exposure-first selection loss efficient mitigate ghost artifacts. Our empirical evaluation validates the effectiveness of the proposed LDR-HDR formation pipeline, as well as experiments show that our method achieves comparable performance to the state-of-the-art methods with less computational cost.